Dissertation title

Vishakh Pradeep Kumar^{1}, Supervisor: Dr Rehan Ahmed

Heriot-Watt University, School of Engineering and Physical Sciences, Mechanical Engineering, Dubai

Abstract

A concise summary of the work and main results in no more than 200 words.

Keywords: 3 to 5 keywords here, in the form: keyword, keyword

1. Introduction

Introduction giving the rationale and background to your work. This also should justify and clearly state your aims and objectives.

There are two typical ways of structuring the introduction. The more compact form is to combine the ‘context’ and ‘literature review’ in a single main section, where the context is presented immediately after the section heading (ie, no subsection), and then structure the literature review as subsection within the first section, finishing off with a final subsection which states the aims and objectives as well as a paragraph describing the structure of the remaining paper.

The more extended form is to separate the context and aims from the literature review. In this case, the first ’Introduction’ section presents the context and concludes with the aims and objective plus the paragraph outlining the structure of the remaining paper. The literature review itself will then be the main Section 2, with suitable subsections to structure the material.

The introduction and literature review usually makes up about 20 to 30% of the overall paper. Make sure that all the material is relevant to your research question and your research. An example of a typical balance of the different sections expected in a standard paper is given in Tab. ??. However, each dissertation is different, and might end up with a different balance, as long as the overall length is not exceeded. In particular, the number of figures, tables and reference can vary widely (and wildly), very much depending on the type of work carried out.

2. Style guide

2.1. Page and text formatting

If you use the provided templates, the style requirements are already the default settings — so don’t tinker with them! This LaTeX template is based on the Elsevier class but using 11pt (instead of the standard 10pt). We use the single-column format for practical reasons.

The document has to be prepared for the UK standard paper of A4 size with a text area of 16.45 cm by 21.9 cm using single columns at a ‘normal’ serif font (e.g., Times New Roman or Cambria) with font size 11pt.

1Student ID : H00428384 Preprint submitted to MSc Advanced Mechanical Engineering July 30, 2025

Table 1: Illustration of relative length of different sections, taken from a representative paper from a previous student.

Section  Word count  Number of figures and tables
Introduction and Literature Review  1500  2 Figures
Apparatus and Methods  1500  3 Figures and 2 Tables
Results  1000 – 1500  5 Figures
Discussion and Conclusions  1000 – 1500  1 Table
Conclusions and Future work  500  1 Table
References  25 – 35 references; 60% journal articles, 10% books, 30% websites
Total  6000  10 Figures and 3 Tables

2.2. Word count

The expected word count is between 5000 and 7000 words. The word count includes everything from the start of the Introduction to the end of the Conclusions, including text in figure captions and tables. Excluded from the word count is the front matter (from the title to the end of the abstracts and key words) and the end matter (acknowledgements, references, appendices).

If you try to cheat the word count by having a lot of important information in appendices: remember that appendices only provide supplementary material, not essential material for the assessment. The markers are not required to read any appendix during the marking of the dissertation.

2.2.1. Section and item numbering

Paragraphs are justified on both sides and start with an indent. Section numbering is numeric, with ‘section’ headings in bold but sub-section and subsub-section headings in italics. Each heading is preceded and followed by some space (about 6pt or half a line).

Figures, tables, and equations are numbered consecutively: Figure 1, Figure 2, Table 1, Table 2, (1), (2), and so on. That means that they are not sub-numbered for each section, so no Figure 1.2. However, a figure might have two or more graphs. In that case, each graph is labelled a), b) and so on. Similarly, equations can be single equations such as

e = mc^{2}  (1)

or they could be a set of equations, using the environment ‘subequation’ from the subcaption package,

C_{p} = \frac{p}{\frac{1}{2}ρU^{2}}  (2a)
C_{P} = \frac{P}{\frac{1}{2}ρAU^{3}}  (2b)

When referring to these objects in the text, you can use either ‘figure ??’, ‘Figure ??’, or ‘Fig ??’, as long as you do it consistently. A specific graph in a multi-graph figure would be referred to as, for example Fig. 2b. Likewise, for referring to a table, you would use table, Table, or Tab. ??, and equations are referred to as Eq. (??), Equations (??) or equation (??).